# On the Self-awareness of Large Reasoning Models’ Capability Boundaries

### v2

@@ 5 Capability Boundary Revealed by Hidden States
The left subfigure of Figure 5 visualizes the classification of solvable and unsolvable questions (see Appendix D for results of all LRMs), along with the capability boundary (blue line) determined by the decision boundary of linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated, under random train – test splits. To quantify this separation, Table 2 shows the classification accuracy, which is consistently close to and robust to the choice of classifier and hyperparameters (e.g., the regularization coefficient of LR). These results demonstrate that the features of capability boundaries are strongly encoded in the hidden states of the last input token.
Interestingly, we also observe that among solvable questions, some lie very close to the capability boundary, while others are far away. We compare these two cases by measuring their token usage during inference. Although LRMs are able to produce correct answers, the average token usage of the questions close to the boundary is – times longer than that far from the boundary (right subfigure of Figure 5). This further supports the validity of the identified capability boundary: solvable questions closer to the boundary require longer reasoning to reach the correct answer.
Therefore, the capability boundary is encoded in the hidden states of the last input token, and it also reflects the reasoning effort required to solve the question.
@@ 6 Reason with Capability Boundary
The analysis in the previous sections reveal that LRMs exhibit clear signs of self-awareness on capability boundaries: reasoning expressions expose confidence patterns that distinguish solvable from unsolvable problems, and hidden states encode boundary information even before reasoning begins. These findings suggest that capability boundaries are not only detectable but can also be harnessed to improve efficiency and reliability of LRMs’ response. In this section, we propose two strategies to reason with capability boundary and demonstrate their effectiveness.
@@ 6.1 Experimental Design
Strategies. Building on the evidence that self-awareness of capability boundaries emerges during test time, we propose two test-time monitoring strategies that leverage boundary-awareness to guide the reasoning process. The key idea is to monitor signals of capability boundaries on the fly, either from reasoning trajectories (black-box) or hidden states (white-box), and to adjust reasoning when questions are beyond the model’s capability. This approach allows LRMs to avoid unproductive reasoning and give more reliable response.
Reasoning Expression Monitoring (Monitor) is a black-box strategy that leverages expressions density during reasoning. The pipeline is as follows:
1.
Construct expression trajectories: continuously track the densities of confident and uncertain expressions (e.g., “I think that’s it” vs. “I might be wrong”) across reasoning steps to form expression density trajectories and .
2.
Detect unsolvable question: analyze these trajectories on the fly using quantitative indicators such as Confidence Differential (ConfDiff) and Confidence Curvature (ConfCurv) to determine whether a question lies within or beyond the capability boundary.
3.
Early stop unproductive reasoning and reprompt with suffix: if the trajectory indicates the question is beyond the capability boundary, we early terminate reasoning and reprompt the model to output a concise approach. This is implemented by appending a self-awareness prompt suffix (shown below) to the end of the question; otherwise, it continues with the standard reasoning process.
1.
@@
1.
Extract hidden states: capture the hidden states of the last input token, which encodes the semantic representation of the entire question during the prefilling stage.
2.
Detect unsolvable question: apply simple linear classifiers (e.g., LDA or LR) trained to distinguish unsolvable from solvable questions. This is performed before reasoning begins.
3.
Avoid unproductive reasoning via constrained output prefix: if the classifier predicts that the question is beyond the capability boundary, the model abstains from full reasoning and instead outputs a concise possible approach. This is implemented by constraining the response to begin with a self-awareness output prefix (shown below); otherwise, the model proceeds with the standard reasoning process.
1.
Notably, both strategies preserve the model’s performance on solvable questions while providing more efficient and reliable responses on questions beyond its capability boundary.
Accuracy (ACC) measures the proportion of correctly solved problems, serving as the standard indicator of reasoning quality (Xia et al., 2025).
Hard Abstention (HA) measures the model’s ability to recognize problems beyond its capability boundary. Different to abstention which focus on underspecification, outdated information, or noncompliance (Kirichenko et al., 2025; Madhusudhan et al., 2024; Brahman et al., 2024), HA specifically evaluates whether LRMs can explicitly abstain on unsolvable problems instead of producing incorrect answers.
Baseline. Prior work has explored improving reasoning performance by self-verification after completing the full Chain-of-Thought (Madhusudhan et al., 2024) or by monitoring undesirable behaviors during training (Baker et al., 2025). However, these approaches are orthogonal to our scenario: the former introduces additional computation after reasoning, while the latter does not address hard questions beyond the capability boundary.
Therefore, we adopt a baseline that improve efficiency by enabling the model to abstain. In particular, we consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of questions beyond capability boundary (see Appendix B for the full prompt).
6.2 Results
@@
Observation 1: our strategies enable LRMs to recognize unsolvable questions without compromising their accuracy on solvable ones.
Although the original models achieve acceptable Accuracy (ACC), their Hard Abstention (HA) is zero, indicating a complete lack of ability to abstain on questions beyond the capability boundary. Moreover, almost all Overflow is 100%, suggesting that most incorrect cases arise because LRMs struggling to solve a hard question until exceeding the context length.
Promisingly, both Monitor and Monitor achieve strong performance. First, their ACC remains nearly unchanged compared to the original models, showing that good reasoning performance is preserved. Meanwhile, the substantially high HA (almost 100%) demonstrates that LRMs can now effectively identify problems beyond the capability boundary.
@@
Observation 2: our strategies enable LRMs to provide more efficient and reliable responses to unsolvable questions by offering concise solution approaches.
Once questions beyond the capability boundary are recognized, our strategies enable LRMs to improve efficiency and reliability by offering concise solution approaches. Figure 6 illustrates the responses produced by our strategies compared to the original models. The original models struggle with unsolvable questions and often fail to complete the reasoning process even after exhausting the context length. Such responses provide neither a correct answer nor valuable insights from the lengthy reasoning trace, while also wasting user time.
In contrast, our strategies detect these unsolvable questions and guide the model to explicitly acknowledge that they are beyond capability boundary, while offering a concise possible approach. This response paradigm provides users with possible hints for problem solving without wasting time on meaningless output. And importantly, it does not compromise the model’s ability to produce correct answers for solvable questions.
This explains why we observe large reductions in both token usage (up to 93.6%) and overflow (up to 100%) in Table 3, indicating that reasoning with capability boundaries greatly improves efficiency and mitigates incomplete outputs. Besides, this reduction becomes more evident with larger context length. For example, overflow reduces from 78.4% to 5.4% on GPT-oss-20B when the context length is extended from 2048 to 4096. This suggests that our strategies holds promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024).
Observation 3: LRMs vary in reasoning behavior correction, with some good at instruction following and others more prone to output guidance.
We compare the performance of Monitor and Monitor across different LRMs. For GPT-oss-20B, Monitor achieves lower token usage than Monitor. We find this model is resistant to output guidance via constrained output prefix: although it sometimes begins by generating a solution approach, it tends to later negate this and restart the reasoning process, resulting in long and unproductive outputs. In contrast, its instruction-following ability is relatively strong, and thus reprompting with the self-awareness prompt suffix yields lower token usage.
Conversely, the other three LRMs show weaker ability of instruction following but more prone to output guidance, making Monitor more effective than Monitor. This suggest that, in addition to the black-box versus white-box distinction, the choice of optimization strategy in practice should also consider the behavioral tendencies of the specific LRM.
@@ 7 Conclusion
In this work, we investigate whether Large Reasoning Models (LRMs) possess self-awareness of their capability boundaries. Through both black-box and white-box analyses, we found consistent evidence that LRMs can distinguish unsolvable from solvable questions: reasoning expressions reveal separable expression trajectories, while hidden states of the last input token encode boundary information even before reasoning begins. Building on these insights, we proposed two simple yet effective test-time monitoring strategies that leverage boundary signals to prevent unproductive reasoning for unsolvable questions. Our experiments show that these strategies preserve accuracy on solvable problems while substantially improving efficiency and reliability on unsolvable ones.
Appendix A Datasets, Models, and Performance
In this section, we provide the specific description of models and datasets evaluated in our work. And we provide the performance of each model on each dataset.
@@ Appendix A Datasets, Models, and Performance
DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025): A 32B-parameter distilled model obtained by transferring reasoning abilities from DeepSeek-R1 into the Qwen2.5-32B backbone, yielding strong performance across multiple benchmarks.
@@
AIME’24 (MAA, 2024): The 2024 American Invitational Mathematics Examination. Similar in format to AIME’25, it consists of 30 problems and provides a prior-year benchmark for evaluating model robustness across exam editions.
@@
AMC’23 (MAA, 2023): The 2023 American Mathematics Competition (AMC 10A). It consists of 30 questions, providing a lighter but diverse benchmark compared to AIME and HMMT.
@@
GSM8K (Cobbe et al., 2021): A large-scale dataset of grade-school math word problems requiring multi-step arithmetic reasoning. We evaluate on approximately 1.5k problems from the test set. It is widely used to evaluate the basic reasoning competence of LRMs, with accuracy typically exceeding 90%.
@@
Humanity’s Last Exam (HLE) (Phan et al., 2025): A multi-modal benchmark at the frontier of human knowledge, designed with broad subject coverage across mathematics, humanities, and natural sciences. We use approximately 1.5k text-only questions (excluding image-based problems) from this dataset. Model accuracy on this dataset is typically below 10%, making it a strong stress test for reasoning beyond capability boundaries.
The performance of LRMs on these Math benchmarks are shown in the following Table 4:
| AIME’25 | AIME’24 | AMC’23 | HMMT | GSM8K | HLE | |
| GPT-oss-20B | ||||||
| R1-Distill-Qwen3-8B | ||||||
| R1-Distill-Qwen-32B | ||||||
| QwQ-32B |
Appendix B System Prompt of BoostAbstention

### v3

@@ 5 Capability Boundary in Hidden States
Since hidden states are high-dimensional (e.g., 4096), limited samples may lead to sparsity issues that obscure statistical patterns (Vishwakarma, 2024; Köppen, 2000). We mix the mathmatical benchmarks used in the previous sections and randomly split it into an 80% training set and a 20% test set, with balanced labels for eventually unsolved and solved questions.
Next, we extract hidden states from the prefilling stage as it processes the input question (Yuan et al., 2024; Pope et al., 2023), using the hook method (Vig, 2019). Specifically, we focus on the hidden state of the last input token, which encapsulates the semantic content of the entire question (Dong et al., 2025; Radford et al., 2019). Finally, we apply simple linear classifiers to these hidden states, including Linear Discriminant Analysis (LDA) and Logistic Regression (LR) (Hastie, 2009). Using final answer correctness as labels, we test whether eventually solved and unsolved questions can be linearly separated, i.e., whether last token hidden states contain information predictive of eventual correctness under our setup.
The left subfigure of Figure 5 visualizes the classification of solvable and unsolvable questions (see Appendix H for results of all LRMs), together with the decision boundary (blue line) of a linear classifier (e.g., LDA). We observe that the two classes of questions are clearly separated under random train–test splits. To quantify this separation, Table 2 reports the classification accuracy, which is consistently close to and robust to the choice of classifier and hyperparameters (e.g., the regularization coefficient of LR). These results suggest that the hidden states of the last input token contain information that is predictive of eventual correctness, and thus closely related to the operational capability boundary under our evaluation setup.
At the same time, these signals should be interpreted with care: they may partially reflect question difficulty rather than a purely intrinsic notion of capability boundary. We provide additional analyses and discussion of this issue in Appendix E.
Interestingly, we also observe that among eventually solved questions, some lie very close to the classifier decision boundary, while others lie farther away. We quantify this distance by the probability difference assigned by the LDA classifier to the solved and unsolved classes (Hwang et al., 2023). We then compare these groups in terms of token usage during inference. Although LRMs answer both groups correctly, questions closer to the boundary require on average – more tokens than those farther from the boundary (right subfigure of Figure 5). This suggests that the hidden state signals are related not only to eventual correctness, but also to the amount of reasoning effort required to obtain a correct answer.
Overall, these results indicate that the hidden states of the last input token contain useful signals related to the operational capability boundary, and that these signals also correlate with the reasoning effort required to solve a question.
Moreover, because these signals are extracted from reasoning expressions and hidden states rather than task-specific heuristics, they may extend beyond mathematical reasoning. We provide an analysis on a coding task in Appendix G.
@@ 6 Reason with Capability Boundary
Because signals related to the operational capability boundary emerge early, they can be used to improve the efficiency and reliability of LRMs’ responses. In this section, we propose two test-time monitoring strategies that leverage such signals and demonstrate their effectiveness.
@@ 6.1 Experimental Design
Strategies. The key idea is to monitor signals of capability boundaries on the fly, either from reasoning trajectories (black-box) or hidden states (white-box), and to adjust reasoning when a question lies beyond the model’s effective solving regime.
Reason Expression Monitoring (Monitor) is a black-box strategy that leverages expressions density during reasoning. The pipeline is:
1. Construct expression trajectories: continuously track the densities of confident and uncertain expressions (e.g., “I think that’s it” vs. “I might be wrong”) across reasoning steps to form expression density trajectories and .
2. Predict likely-unsolved question: analyze these expression trajectories on the fly using quantitative indicators such as Confidence Differential (ConfDiff) to determine whether a question is likely to remain unsolved.
3. Early stop unproductive reasoning and reprompt with suffix: if the trajectory indicates the question is likely to remain unsolved, we early terminate reasoning and reprompt the model to output a concise approach. This is done by appending a prompt suffix at the end of the question:
@@
1. Extract hidden states: capture the hidden states of the last input token, which encodes the semantic representation of the entire question during the prefilling stage.
2. Predict likely-unsolved question: apply simple linear classifiers (e.g., LDA or LR) trained to distinguish eventually unsolved from solved questions. This prediction is made before reasoning begins.
3. Avoid unproductive reasoning via constrained output prefix: if the classifier predicts that the question is likely to remain unsolved under the current setup, the model abstains from full reasoning and instead outputs a concise possible approach. This is implemented by constraining the response to begin with an output prefix (see Appendix F for the ablation study):
@@
Accuracy (ACC) measures the proportion of correct answers, serving as the standard indicator of reasoning quality (Xia et al., 2025).
Hard Abstention (HA) measures the model’s ability to recognize questions beyond its capability boundary. Different to abstention which focus on underspecification, outdated information, or noncompliance (Kirichenko et al., 2025; Madhusudhan et al., 2024; Brahman et al., 2024), HA specifically evaluates whether LRMs can explicitly abstain on unsolvable questions instead of producing incorrect answers.
Baseline. Prior work has explored improving performance by self-verification after generating the answer (Madhusudhan et al., 2024; Kadavath et al., 2022) or by monitoring uncertainty and other undesirable behaviors (Baker et al., 2025; Farquhar et al., 2024; Kuhn et al., 2023). However, these approaches are orthogonal to our scenario: the former introduces additional computation after producing an answer, while the latter does not address eventually unsolved questions (see Appendix B).
Therefore, we adopt a baseline that improve efficiency by enabling the model to abstain. In particular, we consider BoostAbstention (Kirichenko et al., 2025) which crafts a system prompt that encourages the model to abstain in designated scenarios, showing effectiveness for both reasoning and standard LLMs. For our experiments, we modify the system prompt to better match the abstention scenarios of eventually unsolved questions (see Appendix D for the full prompt).
6.2 Results
Observation 1: our strategies enable LRMs to recognize eventually unsolved questions while maintaining comparable accuracy on solved ones. Although the original models achieve acceptable Accuracy (ACC), their Hard Abstention (HA) is zero, indicating a complete lack of ability to abstain on questions beyond the capability boundary. Moreover, almost all Overflow is 100%, suggesting that most incorrect cases arise because LRMs struggling to solve a hard question until context limit.
Promisingly, both strategies achieve strong performance. First, their ACC remains nearly unchanged compared to the original models, showing that good reasoning performance is preserved. Meanwhile, the substantially high HA (almost 100%) demonstrates that LRMs can now effectively identify likely-unsolved questions.
@@ 6.2 Results
Observation 2: our strategies enable LRMs to provide more efficient and reliable responses to eventually unsolved questions by offering concise solution approaches. Figure 6 illustrates the responses produced by our strategies compared to the original models. The original models struggle with eventually unsolved questions and fail to complete the reasoning process even after exhausting the context length. Such responses provide neither a correct answer nor valuable insights from the lengthy reasoning trace, while also wasting time.
In contrast, our strategies detect these unsolved questions and guide the model to explicitly acknowledge that they are beyond capability boundary, while offering a concise possible approach. This response paradigm provides users with possible hints for problem solving without wasting time on meaningless output. And importantly, it does not compromise the model’s ability to produce correct answers for solvable questions.
This explains why we observe large reductions in both token usage (up to 93.6%) and overflow (up to 100%) in Table 3, indicating that reasoning with capability boundaries greatly improves efficiency and mitigates incomplete outputs. Besides, this reduction becomes more evident with larger context length. For example, overflow reduces from 78.4% to 5.4% on GPT-oss-20B when the context length is extended from 2048 to 4096. This suggests that our strategies hold promise for optimizing long-context reasoning tasks (Ling et al., 2025; Kuratov et al., 2024). We provide in Appendix I an additional analysis with the context length set to maximum.
Observation 3: LRMs vary in reasoning behavior correction, with some good at instruction following and others more prone to output guidance. We compare the performance of Monitor and Monitor across different LRMs. For GPT-oss-20B, Monitor achieves lower token usage than Monitor. We find this model is resistant to output guidance via constrained output prefix: although it sometimes begins by generating a solution approach, it tends to later negate this and restart the reasoning process, resulting in long and unproductive outputs. In contrast, its instruction-following ability is relatively strong, and thus reprompting with the prompt suffix via Monitor yields lower token usage.
Conversely, the other three LRMs show weaker ability of instruction following but more prone to output guidance, making Monitor more effective than Monitor. This suggest that, in addition to the black-box versus white-box distinction, the choice of optimization strategy in practice should also consider LRMs’ behavioral tendencies.
@@ 7 Conclusion
We study whether signals related to the operational capability boundary can help mitigate unproductive reasoning in LRMs. Across both black-box and white-box analyses, we find that LRMs expose such signals early: reasoning expressions differ systematically between eventually solved and unsolved questions, while the hidden states of the last input token are predictive of eventual correctness even before reasoning begins. Based on these findings, we propose two test-time monitoring strategies that reduce unproductive reasoning. Experiments show that they largely preserve accuracy on solvable questions while substantially improving efficiency and reliability on questions lie beyond the operational capability boundary.
Limitations
Interpretation of capability boundaries. In this paper, we use capability boundary in an operational sense, namely as the empirical separation between questions a model eventually solves and those it does not under a fixed evaluation setup. Accordingly, the signals we identify should be interpreted as predictive correlates of eventual failure rather than definitive evidence of a formally defined intrinsic capability boundary. In particular, these signals may partially reflect question difficulty. Although we attempt to mitigate this concern through per-benchmark and cross-benchmark analyses (Appendix E), the extent to which these signals capture a general notion of capability boundary remains an open question. Developing a more principled formalization of capability boundaries (e.g., disentangling question difficulty) is an important direction for future work.
Scope of tasks and evaluation domains. Our evaluation focuses primarily on mathematical benchmarks, which allow systematic analysis of unsolvability signals but may not fully represent other reasoning scenarios. While our experiments on coding tasks suggest similar results (Appendix G), the generality of capability boundary signals across broader domains (e.g., complex decision making or scientific reasoning) remains to be established. Future work should investigate whether similar strategies apply across more diverse tasks.
Ethics Statement
ACL Ethics Policy is respected in this work. It studies LRMs’ capability boundaries, aiming to improve efficiency and reliability. No human subjects, sensitive personal data, or harmful content were involved. The proposed methods are intended for research purposes only and do not pose foreseeable risks of misuse.
Appendix A Datasets and Models
In this section, we provide the specific description of models and datasets evaluated in our work.
@@ Appendix A Datasets and Models
DeepSeek-R1-Distill-Qwen-32B (Guo et al., 2025): A 32B-parameter distilled model obtained by transferring reasoning abilities from DeepSeek-R1 into the Qwen2.5-32B backbone, yielding strong performance across benchmarks.
@@
For the sampling settings, we use greedy decoding to ensure reproducibility of results. For the maximum sequence length, we follow the specifications in the HuggingFace model cards: 128k for GPT-oss-20B, 64k for DeepSeek-R1-0528-Qwen3-8B, and 32k for both DeepSeek-R1-Distill-Qwen-32B and QwQ-32B.
@@
AIME’24 (MAA, 2024): The 2024 American Invitational Mathematics Examination. Similar in format to AIME’25, it consists of 30 problems and provides a prior-year benchmark for evaluating model robustness across editions.
@@
AMC’23 (MAA, 2023): The 2023 American Mathematics Competition (AMC 10A). The dataset consists of 30 questions, providing a lighter but diverse benchmark compared to AIME and HMMT.
@@
GSM8K (Cobbe et al., 2021): A large-scale dataset of grade-school math word problems requiring multi-step arithmetic reasoning. We evaluate on approximately 1.5k problems from the test set. It is widely used to evaluate the basic reasoning competence of LRMs.
@@
Humanity’s Last Exam (HLE) (Phan et al., 2025): A multi-modal benchmark at the frontier of human knowledge, designed with broad subject coverage across mathematics, humanities, and natural sciences. We use approximately 1.5k text-only questions (excluding image-based problems) from this dataset.
Appendix B Limitations of Some Prior Work
Some prior works evaluate model confidence through self-verification (Kadavath et al., 2022) or uncertainty estimation (Farquhar et al., 2024; Kuhn et al., 2023). In this section, we show that these approaches are less effective than our proposed monitoring strategies for identifying unsolvable questions with comparable accuracy on solvable ones. We first clarify how we adapt these methods to fit our task.
Self-verification (SV). For each question, we ask the model to evaluate whether its own predicted answer is correct using the prompt template:
This follows prior work (Kadavath et al., 2022) which study whether language models can evaluate the validity of their own claims by asking models to first propose answers, and then to evaluate that their answers are correct. If the model judges its own incorrect answer to be incorrect, we count this as a hard abstention (HA). Conversely, if the model judges its correct answer to be incorrect, this reduces the overall task accuracy (Acc). Formally,
| Model judges the answer to be correct | (4) | |||
| Model judges the answer to be incorrect | (5) | |||
Uncertainty-estimation (UE). For each question, we compute the semantic entropy of the model’s generated answer, where a higher entropy indicates greater uncertainty. This follows prior work (Farquhar et al., 2024; Kuhn et al., 2023), which introduces semantic entropy as an uncertainty measure that captures linguistic invariances induced by shared meanings. When the entropy exceeds a certain threshold, the model is considered uncertain about its answer. Similar as before, we define:
| Method | ACC (%) ↑ | HA (%) ↑ |
| GPT-oss-20B | 74.6 | 0 |
| +SV | 32.1 | 46.0 |
| +UE | 50.0 | 29.0 |
| +Monitor | 74.6 | 100 |
| +Monitor | 73.7 | 98.9 |
Results. We provide the results of these two methods alongside our Monitor and Monitor in Table 4. Although both methods provide models with a certain degree of hard abstention capability, they also incorrectly classify many solvable questions as unsolvable, which significantly reduces overall accuracy. In contrast, our approaches enable LRMs to reliably identify unsolvable questions without compromising accuracy on solvable ones.
Appendix C Confident and Uncertain Expressions
To ensure that our analysis is not tied to particular LRM’s reasoning style, we extract confident and uncertain expressions individually for each model, based on its own characteristic reasoning patterns. The extraction pipeline is outlined below.
Step 1: Identify all anthropomorphic expressions from the model’s reasoning trace. Prior work (Guo et al., 2025; Qian et al., 2025) show that the anthropomorphic expressions (e.g., “wait”) are crucial for LRM reasoning performance, in contrast to purely neutral mathematical steps (e.g., “11 + 5 = 16”). Following this insight, we curate a set of anthropomorphic expressions that the LRM commonly uses. And we further verify with GPT-5 that these expressions indeed convey anthropomorphic tone rather than purely propositional content.
Step 2: Select expressions whose occurrence frequency is most predictive of answer correctness. For each question, we count the occurrences of all extracted expressions within reasoning. Using the final answer correctness as the label and the occurrence counts for each expressions as features, we apply feature-importance analysis methods (Wang et al., 2024), including SHAP and Permutation Importance (PI), to identify the subset of expressions most informative for predicting whether the model’s final answer is correct. As shown in Table 5 (an example for GPT-oss-20B), we select expressions such as “wait”, “yes”, and “exactly” which rank among the most predictive, whereas others (e.g., “so”) exhibit negligible predictive value and are therefore excluded.
| Expression | SHAP | PI | Total |
| wait | 1.00 | 1.00 | 2.00 |
| yes | 0.33 | 0.73 | 1.06 |
| exactly | 0.26 | 0.73 | 0.99 |
| so | 0.01 | 0.00 | 0.01 |
Step 3: Divide the selected expressions into confidence and uncertainty categories. This step is grounded in the connection between epistemic markers and expressed uncertainty. Prior work (Liu et al., 2025; Yona et al., 2024) show that epistemic markers reflect LLMs’ internal confidence, analogous to how humans use expressions such as “I am not sure” or “it is unlikely that” to convey uncertainty. Following this insight, we classify each selected expression as reflecting confidence or uncertainty (further validated by GPT-5).
Figure 7 shows the full list of expressions of all LRMs. When matching these expressions, the Uncertain Expression is matched first, followed by the Confident Expression. Additionally, both the position of the first token of an expression and the token length of the entire expression are considered.
We estimate the model’s confidence level from the occurrences of confidence and uncertainty expressions throughout the reasoning process, following prior work (Liu et al., 2025) which measures LLM uncertainty by counting the frequency of epistemic markers. In Section 3, we use the category of expressions that appears more frequently near the end of the reasoning trace to determine whether the model is more confident or more uncertain. Building on this idea, Section 4 analyzes the trajectory of confidence and uncertainty signals over the entire reasoning sequence, captured by ConfDiff (Equation 2) and ConfCurv (Equation 3), to assess whether the model is more confident or more uncertain.
Appendix D System Prompt of BoostAbstention